# Control of kinetic opinion dynamics in popularity-adaptive social networks

### v1

@@ 1 Introduction
Kinetic theory has become a central mathematical framework for the description of large systems of interacting agents, providing tractable models capable of capturing the emergence of collective behaviours across a wide spectrum of applications, ranging from economics and traffic dynamics to the social and biological sciences [PT13, Bellomo]. In these settings, Boltzmann-type equations allow to derive mesoscopic descriptions starting from simple interaction rules, thus linking microscopic decision processes to macroscopic observables. Applications of such paradigm span socio–economic models for price formation and wealth redistribution [Cor05, Par14, Cha07, T22, Burger, DPT18], kinetic descriptions of population and epidemic dynamics [Dim21, Cha04, BDP21], and applications in transportation engineering and crowd modelling [albi2021optimized, cristiani2014multiscale, Gun02, Pup16, ABFH19, BGQR22, Wolfram].
@@
Within this kinetic framework, opinion formation is a central example for interaction mechanisms, where repeated social exchanges give rise to dynamics well described by Boltzmann-type models, capturing consensus, polarisation and fragmentation as studied in [To06, Gal82, DT19, Dim20, fakenews]. Of particular importance is the extension of these modelling frameworks to account for an additional layer of complexity such as individual popularity, typically quantified by the number of connections or followers in a social networks. Hence, recent developments have focused on coupling opinion evolution with the dynamics of the underlying social connectivity structure, see for example [Ace11, Albi24, quattrociocchi2014opinion, fagioli2024opinion].
@@
Here we focus on kinetic model for the evolution of joint density of opinions and contacts, where each agent is represented by an opinion $v\in[-1,1]$ and a popularity variable $c\geq 0$ measuring its social connections. Following the approach proposed in [albi2017opinion, Albi24], we assume that dynamics of opinions depend on agents’ popularity. Furthermore, we introduce a feedback mechanism in which agents’ opinions influence the evolution of their social connections, while the strength of opinion interactions remains modulated by the individuals’ relative popularity. This second assumption has not been considered in previous works within this context, nevertheless it is supported by empirical evidence showing that individuals tend to form or dissolve connections based on opinion similarity. This phenomenon is widely documented in studies on homophily, echo chambers, and opinion-driven rewiring in online social networks, see for instance [cinelli2021echo, himelboim2013tweeting, mcpherson2001birds, nguyen2020echo].
@@
In this modelling framework, we also investigate the impact of control mechanisms acting on both the contact dynamics and the opinion evolution. Control strategies in opinion modelling have been explored at various scales see for example [during2024breaking, albi2025data, bicego2025computation]. Here we follow the approach introduced in [APZ14, albi2015kinetic], deriving two type of feedback controls, the first acting over the contact formation and designed to counteract the endogenous decay of contacts by mitigating unfavourable trends in the visibility of selected agents or groups. A second control term acts at the opinion level, steering agents toward prescribed target positions whenever activation criteria based on connectivity or local opinion density are satisfied.
@@
The integration of such kinetic framework with control thus offers a rigorous tool for designing strategies that promote consensus, limit fragmentation, or may regulate the impact of highly connected agents in public discourse. Overall, the finding of the present manuscript provides a modelling and computational setting for to analyise interventions in complex social systems where popularity, influence, and opinions co-evolve.
@@
The paper is structured as follows. In Section 2, starting from a system of indistinguishable agents, we present the microscopic model for the evolution of the contacts. Such evolution is driven both by the desire of the agents to increase their followers once they enter the network and, consequently, gain more influence, but it also depends on how popular and endorsed are their opinions. In Section 2.1 we introduce the control variable for the evolution of the contacts and the optimal control problem linked to it. The objective of this control variable is to minimize the relative loss of the agent’s contacts. We also present an interpretation of the optimal control functional as the continuous time limit of a discrete-time formulation. We conclude this section with the mean field description of the contacts evolution obtained through a suitable scaling of the contacts upgrade and the frequency of such evolution. Section 3 is devoted to the controlled evolution of the opinions. Again, we start from the microscopic multi-agent description of the binary interactions that shape each agent’s opinion and present the optimal control problem linked to them. As we did for the evolution of the contacts, also for the opinion evolution we present the mean-field limit, leading to the Fokker-Planck-type equation for the time evolution of the joint density of opinions and contacts in the regime of small updates of the opinions and the contacts happening with a very high frequency. Finally, Section 4 presents a set of numerical experiments carried on using a Monte Carlo approach, that show the different outcomes of the joint contacts and opinion evolution under various possible configurations.
@@ 2 The evolutionary model for social network contacts
In this section we introduce the kinetic model governing the evolution of social media contacts. We first assume that agents’ opinions are fixed, and we specify the microscopic interaction rules governing the evolution of each agent’s number of connections. Following the approaches proposed in [Albi24, Dim20], these rules lead to a Boltzmann-type master equation describing the evolution of the corresponding distribution.
@@
$\int_{[-1,1]\times\mathbb{R}_{+}}f(v,c,t)\mathrm{d}v\mathrm{d}c=1,$
@@
$h(c,t):=\int_{[-1,1]}f(v,c,t)\mathrm{d}c,\qquad g(v,t):=\int_{\mathbb{R}_{+}}f(v,c,t)\mathrm{d}v.$
We first analyze the evolution of agents’ connection, considering the opinions as a fixed parameter. Thus, for any given measurable subset $D\subseteq\mathbb{R}_{+}$, the quantity
@@
As in [Albi24], we assume that the agents are generally inclined to increase their number of social media followers once they enter the network, and that they tend to reach, at least, a certain number of followers $\overline{c}$, which represents a desirable popularity level. In addition, we assume that an agent’s ability to acquire or retain contacts is influenced by the popularity of the opinion they express, so that opinion–driven visibility plays a direct role in the dynamics of the contact variable.
@@
$c^{\prime}=c-\beta\left(\Psi\left(\frac{c}{\overline{c}}\right)+\Phi(v)\right)c+\eta c,$ (1)
@@
$\Psi(s)=\frac{1}{\beta}\left(\frac{\mu}{1-\mu}\right)\frac{s-1}{\frac{1+\mu}{1-\mu}s+1},$
with $s=c/\overline{c}$. This function is dimensionless, increasing and it is equal to zero at the reference point $s=1$. Moreover, it is asymmetric: moving toward the reference level $s=1$ is easier for agents starting below the threshold, while agents above the reference level face a weaker incentive to reduce their connections. This reflects the empirical principle that it is easier to gain new contacts when starting from a low number, whereas highly connected users are not motivated to decrease their popularity. This mechanism encodes the principles of the prospect theory formulated by Kahneman and Tversky [KTa]. The function is bounded by the values
$\frac{1}{\beta}\frac{\mu}{1-\mu}\leq\Psi(s)\leq\frac{1}{\beta}\frac{\mu}{1+\mu},$
and it can be seen represented in Figure 1. This specific choice is justified by the results presented in [Albi24].
Figure 1: Plot of $\Psi(s)$. The red lines represent the bounds of the function when $\beta=1$ and $\mu=0.1$.
@@
where $\theta>0$, $m_{v}$ is the mean opinion of the entire population and $\delta_{\Phi}$ is a fixed parameter that depends on the tolerance of the population to unpopular opinions. This function is negative if the agent has an opinion with a distance which is less or equal than $\delta_{\Phi}$ from the average opinion of the population, positive , and due to the compactness of the domain it is bounded in
@@ 2.1 The controlled evolution of social media contacts
We now introduce an external intervention mechanism aimed at regulating the evolution of an agent’s popularity on the social network. To this aim, we modified the uncontrolled dynamics in (1) adding a control input $\kappa$ so that the microscopic evolution becomes
$c^{\prime}=c-\beta\left(\Psi\left(\frac{c}{\overline{c}}\right)+\Phi\left(v\right)-\kappa\right)c+\eta c.$ (2)
The optimal control $\kappa^{\star}$ is obtained as the minimizer of the following cost functional
$\kappa^{\star}=\arg\min_{\kappa}\,\mathcal{J}[\kappa],\qquad\mathcal{J}[\kappa]:=\mathbb{E}\!\left[-\lambda\,\frac{c^{\prime}-c}{c}\,R_{c}(c)\,H_{c}(v)+\frac{\gamma_{c}}{2}\,\kappa^{2}\right],$ (3)
which penalizes the relative loss of the agent’s contacts along with a quadratic regularization term. Indeed, the quantity ${(c^{\prime}-c)}/{c}$ is the incremental relative variation of contacts between two consecutive time steps, and overall the first term of $\mathcal{J}$ promotes control actions capable of counterbalancing an expected loss of contacts, but only in the event that certain activation conditions are met. Specifically, such conditions are expressed by the activation functions $R_{c}(c)$ and $H_{c}(v)$. The parameter $\lambda$ sets the intrinsic scale of the control action, determining the overall intensity with which contact growth can be enhanced relative to the natural dynamics. The second term appearing in the functional $\mathcal{J}$, meaning ${\gamma_{c}\kappa^{2}}/{2}$, penalizes the control magnitude, with the positive constant $\gamma_{c}$ determining how costly it is to increase the number of contacts. The activation function $R_{c}(c)$ only depends on the current number of contacts of the agent and is modeled by the sigmoid
$R_{c}(c)=\frac{1}{1+\exp(-\alpha_{R}(c_{\min}-c))},$
which is close to one when contacts fall below the threshold $c_{\min}$, and decreases as $c$ grows beyond it. The activation function $H_{c}(v)$ depends instead on the local opinion density $m(v)$, it ensures that the control only acts on leaders positioned within a prescribed opinion region and is given by
$H_{c}(v)=\frac{1}{1+\exp(-\alpha_{H}(\rho(v)-\rho^{\star}))},$
@@
$\rho(v)=\int_{\lvert w-v\rvert\leq r}f(w,c,t)\mathrm{d}w\mathrm{d}c,$
i.e. the fraction of agents whose opinions have a distance from $v$ which is less or equal than $r>0$ in, while $\rho^{\star}$ indicates an activation threshold, meaning the minimum mass required for a significant group effect to be witnessed. The parameters $\alpha_{R},\alpha_{H}>0$ modify the steepness on the sigmoid functions. In this framework, $H_{c}(v)$ increases with the density of similar opinions, modeling the social tendency of the agents to seek more contacts when they feel part of a group with a shared vision.
@@
In order to obtain an explicit expression of the optimal control, for a single agent, we follow a greedy approach as in [Albi16], looking for a low complexity computational representation of the optimal control problem expressed in (2)-(3) using the so-called model predictive control, that computes a suboptimal solution by iterating on a sequence of finite time steps. Here, we limit ourselves to a single evolution step, from $c$ to $c^{\prime}$, because we are interested in an instantaneous control strategies. For each agent, we can obtain an explicit expression of the optimal control deriving the functional $\mathcal{J}[\kappa]$ with respect to $\kappa$ and setting the derivative equal to zero. Noticing that $\partial_{\kappa}c^{\prime}=\beta c$, we get
$\partial_{\kappa}\mathcal{J}[\kappa]=\mathbb{E}\!\left[-\lambda\beta R_{c}(c)H_{c}(v)+\gamma_{c}\kappa\right].$
Imposing the first order optimality conditions $\partial_{\kappa}\mathcal{J}[\kappa]=0$, recalling that $\mathbb{E}\left[\eta\right]=0$, and using that $\gamma_{c}>0$ we get
$\lambda\beta R_{c}(c)H_{c}(v)+\gamma_{c}\kappa=0.$
@@
$\kappa^{\star}=\lambda\,\frac{\beta}{\gamma_{c}}\,R_{c}(c)\,H_{c}(v).$ (4)
Remark 2.1.
(Time discrete functional and continuous limit). We notice that the functional (3) can be interpreted as a one-step discrete-time functional. Indeed, introducing a partition of the time interval $[0,T]$, given by $t_{n}=n\Delta t$, $n=0,\dots,N$, $N\Delta t=T$, with $\Delta t>0$, we consider analogously with (3) the following discrete cost functional
$\mathcal{J}_{\Delta t}\;=\;\sum_{n=0}^{N-1}\Delta t\;\mathcal{L}_{n},\qquad\mathcal{L}_{n}\;=\;-\lambda\,\frac{c^{n+1}-c^{n}}{\Delta t\,c^{n}}\,R_{c}(c^{n})H_{c}(v^{n})\;+\;\frac{\gamma_{c}}{2\Delta t}\,(\kappa^{n})^{2}.$
@@
$\mathcal{J}[\kappa]\;=\;\int_{0}^{T}\Big[-\lambda\,\frac{\dot{c}(t)}{c(t)}\,R_{c}(c(t))H_{c}(v(t))\;+\;\frac{\tilde{\gamma}_{c}}{2}\,\kappa(t)^{2}\Big]\,\mathrm{d}t.$ (5)
@@
In order to pass to a mean field description of contacts, we suppose that the process of network formation is the result of small increments of the number of connections of the agents. This translates in the scaling of equation (2), taking Let $\varepsilon>0$, the scaled equation now reads
$c^{\prime}=c-\varepsilon\beta\left(\Psi^{\varepsilon}\left(\frac{c}{\overline{c}}\right)+\Phi(v)-\kappa\right)c+\eta_{\varepsilon}c$
where $\eta_{\varepsilon}$ is a random variable with mean equal to zero and standard deviation $\sqrt{\varepsilon}\nu$, while the value function $\Psi^{\varepsilon}$ is defined as
$\Psi^{\varepsilon}(s)=\frac{1}{\beta\varepsilon}\left(\frac{\mu}{1-\mu}\right)\frac{s^{\varepsilon}-1}{\frac{1+\mu}{1-\mu}s^{\varepsilon}+1},$
@@
Supposing $v$ fixed, we can now derive an evolutionary model for the controlled formation of the network. The rescaled density of the connections, $h_{\varepsilon}(v,c)$, obeys a Boltzmann-like equation of the form
$\partial_{t}\int_{\mathbb{R}_{+}}\varphi(c)h_{\varepsilon}(v,c)\mathrm{d}c=\mathbb{E}\left[\int_{\mathbb{R}_{+}}\chi(c)\left(\varphi(c^{\prime})-\varphi(c)\right)h_{\varepsilon}(v,c)\mathrm{d}c\right],$ (6)
@@
$\mathbb{E}\left[c^{\prime}-c\right]=\varepsilon\beta\left(\Psi^{\varepsilon}\left(\frac{c}{\overline{c}}\right)+\Phi(v)-\kappa\right)c,$ (7)
@@
$\mathbb{E}\left[(c^{\prime}-c)^{2}\right]=\varepsilon^{2}\beta^{2}\left(\Psi^{\varepsilon}\left(\frac{c}{\overline{c}}\right)+\Phi(v)-\kappa\right)^{2}c^{2}+\varepsilon\nu^{2}c^{2},$ (8)
@@
$\mathbb{E}\left[(c^{\prime}-c)^{3}\right]=$ $\varepsilon^{3}\beta^{3}\left(\Psi^{\varepsilon}\left(\frac{c}{\overline{c}}\right)+\Phi(v)-\kappa\right)^{3}c^{3}+\mathbb{E}[\eta_{\varepsilon}^{3}]c^{3}$ (9)
$+3\varepsilon^{2}\beta\nu^{2}\left(\Psi^{\varepsilon}\left(\frac{c}{\overline{c}}\right)+\Phi(v)-\kappa\right)c^{3}.$

### v2

@@ 1 Introduction
Kinetic theory has become a central mathematical framework for the description of large systems of interacting agents, providing tractable models capable of capturing the emergence of collective behaviors across a wide spectrum of applications, ranging from economics and traffic dynamics to the social and biological sciences [34, 10]. In these settings, Boltzmann-type equations allow to derive mesoscopic descriptions starting from simple interaction rules, thus linking microscopic decision processes to macroscopic observables. Applications of such paradigm span socio–economic models for price formation and wealth redistribution [18, 35, 16, 39, 14, 24], kinetic descriptions of population and epidemic dynamics [20, 15, 13], and applications in transportation engineering and crowd modelling [6, 19, 29, 36, 2, 11, 26].
@@
Within this kinetic framework, opinion formation is a central example for interaction mechanisms, where repeated social exchanges give rise to dynamics well described by Boltzmann-type models, capturing consensus, polarisation and fragmentation as studied in [38, 28, 21, 22, 27]. Of particular importance is the extension of these modelling frameworks to account for an additional layer of complexity such as individual popularity, typically quantified by the number of connections or followers in a social networks. Hence, recent developments have focused on coupling opinion evolution with the dynamics of the underlying social connectivity structure, see for example [1, 5, 37, 25].
@@
Here we focus on kinetic model for the evolution of joint density of opinions and contacts, where each agent is represented by an opinion $v\in[-1,1]$ and a popularity variable $c\geq 0$ measuring its social connections. Following the approach proposed in [9, 5], we assume that dynamics of opinions depend on agents’ popularity. Furthermore, we introduce a feedback mechanism in which agents’ opinions influence the evolution of their social connections, while the strength of opinion interactions remains modulated by the individuals’ relative popularity. This second assumption has not been considered in previous works within this context, nevertheless it is supported by empirical evidence showing that individuals tend to form or dissolve connections based on opinion similarity. This phenomenon is widely documented in studies on homophily, echo chambers, and opinion-driven rewiring in online social networks, see for instance [17, 30, 32, 33].
@@
In this modelling framework, we also investigate the impact of control mechanisms acting on both the contact dynamics and the opinion evolution. Control strategies in opinion modelling have been explored at various scales see for example [23, 3, 12]. Here we follow the approach introduced in [8, 7], deriving two type of feedback controls, the first acting over the contact formation and designed to counteract the endogenous decay of contacts by mitigating unfavourable trends in the visibility of selected agents or groups. A second control term acts at the opinion level, steering agents toward prescribed target positions whenever activation criteria based on connectivity or local opinion density are satisfied.
@@
The integration of such kinetic framework with control thus offers a rigorous tool for designing strategies that promote consensus, limit fragmentation, or may regulate the impact of highly connected agents in public discourse. Overall, the finding of the present manuscript provides a modeling and computational setting for to analyze interventions in complex social systems where popularity, influence, and opinions co-evolve.
@@
The paper is structured as follows. In Section 2, starting from a system of indistinguishable agents, we present the microscopic model for the evolution of the contacts. Such evolution is driven both by the desire of the agents to increase their followers once they enter the network and, consequently, gain more influence, but it also depends on how popular and endorsed are their opinions. In Section 2.1 we divide the agents in different populations and we introduce the control variable for the evolution of the contacts and the optimal control problem linked to it. The objective of this control variable is to minimize the relative loss of the agent’s contacts. We also present an interpretation of the optimal control functional as the continuous time limit of a discrete-time formulation. We conclude this section with the mean field description of the contacts evolution obtained through a suitable scaling of the contacts upgrade and the frequency of such evolution. Section 3 is devoted to the controlled evolution of the opinions. Again, we divide the agents in distinct populations, then we start from the microscopic multi-agent description of the binary interactions that shape each agent’s opinion and present the optimal control problem linked to them. As we did for the evolution of the contacts, also for the opinion evolution we present the mean-field limit, leading to the Fokker-Planck-type equation for the time evolution of the joint density of opinions and contacts in the regime of small updates of the opinions and the contacts happening with a very high frequency. Finally, Section 4 presents a set of numerical experiments carried on using a Monte Carlo approach, that show the different outcomes of the joint contacts and opinion evolution under various possible configurations.
@@ 2 The evolutionary model for social network contacts
In this section we introduce the kinetic model governing the evolution of social media contacts. We first assume that agents’ opinions are fixed, and we specify the microscopic interaction rules governing the evolution of each agent’s number of connections. Following the approaches proposed in [5, 22], these rules lead to a Boltzmann-type master equation describing the evolution of the corresponding distribution.
@@
$\iint_{[-1,1]\times\mathbb{R}_{+}}f(v,c,t)\mathrm{d}v\mathrm{d}c=1,$
@@
[…]
We first analyze the evolution of agents’ connections, considering the opinions as a fixed parameter. Thus, for any given measurable subset $D\subseteq\mathbb{R}_{+}$, the quantity
@@
As in [5], we assume that the agents are generally inclined to increase their number of social media followers once they enter the network, and that they tend to reach, at least, a certain number of followers $\overline{c}$, which represents a desirable popularity level. In addition, we assume that an agent’s ability to acquire or retain contacts is influenced by the popularity of the opinion they express, so that opinion–driven visibility plays a direct role in the dynamics of the contact variable.
@@
$\color[rgb]{0,0,0}{c^{\prime}=c-\left(\Psi\left(\frac{c}{\overline{c}}\right)+\beta\Phi(v)\right)c+\eta c}$}, (2)
@@
$\color[rgb]{0,0,0}{\Psi(s)=\left(\frac{\mu}{1-\mu}\right)\frac{s-1}{\frac{1+\mu}{1-\mu}s+1}}$},
with $s=c/\overline{c}$. This function is dimensionless, increasing and it is equal to zero at the reference point $s=1$. Moreover, it is asymmetric: moving toward the reference level $s=1$ is easier for agents starting below the threshold, while agents above the reference level face a weaker incentive to reduce their connections. This reflects the empirical principle that it is easier to gain new contacts when starting from a low number, whereas highly connected users are not motivated to decrease their popularity. This mechanism encodes the principles of the prospect theory formulated by Kahneman and Tversky [31]. The function is bounded by the values
$\color[rgb]{0,0,0}{-\frac{\mu}{1-\mu}\leq\Psi(s)\leq\frac{\mu}{1+\mu}}$},
and it can be seen represented in Figure 1. This specific choice is justified by the results presented in [5].
Figure 1: Plot of $\Psi(s)$. The red lines represent the bounds of the function when $\mu=0.1$.
@@
where $\theta>0$, $m_{v}$ is the mean opinion of the entire population and $\delta_{\Phi}$ is a fixed parameter that depends on the tolerance of the population to unpopular opinions. This function is negative if the agent has an opinion with a distance which is less or equal than $\delta_{\Phi}$ from the average opinion of the population, positive otherwise, and due to the compactness of the domain it is bounded in
@@ 2.1 The controlled evolution of social media contacts
In the following, we assume that the agents are divided into a finite number of distinct populations, each of them possibly characterized by a different control strategy acting on the evolution of contacts and/or opinions. We denote by $f_{\ell}(v,c,t)$ the density of agents belonging to the $\ell$-th population with opinion $v\in[-1,1]$ and number of contacts $c\in\mathbb{R}_{+}$ at time $t$. The total density is then given by
$f(v,c,t):=\sum_{\ell}f_{\ell}(v,c,t),$
and the marginal densities of opinions and contacts are defined analogously to (1). Accordingly, the mass of each population is defined as
$\rho_{\ell}:=\iint_{[-1,1]\times\mathbb{R}_{+}}f_{\ell}(v,c,t)\,dv\,dc,\,\textrm{ with }\sum_{\ell}\rho_{\ell}=1,$
where each population is conserved in time, since there is no exchange of agents between different populations, which can be interpreted as a compartment characterized by its own control strategy. The mean contacts and opinions write respectively as follows
$\rho_{\ell}m_{\ell,c}(t)$ $=\iint_{[-1,1]\times\mathbb{R}_{+}}cf_{\ell}(v,c,t)\,dv\,dc,$
$\rho_{\ell}m_{\ell,v}(t)$ $=\iint_{[-1,1]\times\mathbb{R}_{+}}vf_{\ell}(v,c,t)\,dv\,dc.$
We now describe the evolution of the agent’s popularity of population $\ell$, modifying the uncontrolled dynamics in (2) introducing a control input $\kappa_{\ell}$ in such dynamics:
$\color[rgb]{0,0,0}{c^{\prime}=c-\left(\Psi\left(\frac{c}{\overline{c}}\right)+\beta\left(\Phi\left(v\right)-\kappa_{\ell}\right)\right)c+\eta c.}$} (3)
In this work, we restrict ourselves to considering a common functional form of the contact control for all populations, and we distinguish the different groups only through the values of the corresponding control parameters. This choice allows us to model heterogeneous intervention capabilities within a unified control framework, while keeping the analytical structure of the model unchanged across populations.
The optimal control $\kappa_{\ell}^{\star}$ is obtained as the minimizer of the following cost functional
[…]
which penalizes the relative loss of the agent’s contacts along with a quadratic regularization term. Indeed, the quantity ${(c^{\prime}-c)}/{c}$ is the incremental relative variation of contacts between two consecutive time steps, and overall the first term of $\mathcal{J}_{\ell}$ promotes control actions capable of counterbalancing an expected loss of contacts, but only in the event that certain activation conditions are met. Specifically, such conditions are expressed by the activation functions $R^{\ell}_{c}(c)$ and $H^{\ell}_{c}(v)$. The parameter $\lambda^{\ell}$ sets the intrinsic scale of the control action, determining the overall intensity with which contact growth can be enhanced relative to the natural dynamics. The second term appearing in (4), penalizes the control magnitude, with the positive constant $\gamma^{\ell}_{c}$ determining how costly it is to increase the number of contacts. The activation function $R^{\ell}_{c}(c)$ only depends on the current number of contacts of the agent and is modeled by the sigmoid
$R^{\ell}_{c}(c)=\frac{1}{1+\exp(-\alpha^{\ell}_{R}(c^{\ell}_{\min}-c))},$
[…]
$H^{\ell}_{c}(v)=\frac{1}{1+\exp(-\alpha^{\ell}_{H}(\rho(v)-\rho^{\ell}_{\star}))},$
@@
$\rho(v)=\iint_{\lvert w-v\rvert\leq r}f(w,c,t)\mathrm{d}w\mathrm{d}c,$
i.e. the fraction of agents whose opinions have a distance from $v$ which is less or equal than $r>0$, while $\rho^{\ell}_{\star}$ indicates an activation threshold, meaning the minimum mass required for a significant group effect to be witnessed. The parameters $\alpha^{\ell}_{R},\alpha^{\ell}_{H}>0$ modify the steepness on the sigmoid functions. In this framework, $H^{\ell}_{c}(v)$ increases with the density of similar opinions, modeling the social tendency of the agents to seek more contacts when they feel part of a group with a shared vision.
@@
In order to obtain an explicit expression of the optimal control for a single agent, we follow a greedy approach as in [4], looking for a low complexity computational representation of the optimal control problem expressed in (3)-(4) using the so-called model predictive control, that computes a suboptimal solution by iterating on a sequence of finite time steps. Here, we limit ourselves to a single evolution step, from $c$ to $c^{\prime}$, because we are interested in an instantaneous control strategy. For each agent, we can obtain an explicit expression of the optimal control deriving the functional $\mathcal{J}_{\ell}[\kappa]$ with respect to $\kappa$ and setting the derivative equal to zero. Indeed, noticing that the expected value of $\eta$ is equal to zero, and that $\partial_{\kappa}c^{\prime}=\beta c$, we get
[…]
Imposing the first order optimality conditions $\partial_{\kappa}\mathcal{J}_{\ell}[\kappa]=0$ and using that $\gamma^{\ell}_{c}>0$ we get
$\lambda^{\ell}R^{\ell}_{c}(c)H^{\ell}_{c}(v)+\gamma^{\ell}_{c}\kappa=0.$
@@
$\kappa^{\star}_{\ell}=\frac{\lambda^{\ell}}{\gamma_{c}^{\ell}}\,R^{\ell}_{c}(c)\,H^{\ell}_{c}(v).$ (5)
Remark 2.1 (Time discrete functional and continuous limit.).
We notice that the functional (4) can be interpreted as a one-step discrete-time functional. Indeed, introducing a partition of the time interval $[0,T]$, given by $t_{n}=n\Delta t$, $n=0,\dots,N$, $N\Delta t=T$, with $\Delta t>0$, we consider analogously with (4) the following discrete cost functional (for the sake of notational simplicity, we omit here the population index $\ell$):
$\mathcal{J}_{\Delta t}\;=\;\sum_{n=0}^{N-1}\Delta t\;\mathcal{L}_{n},\qquad\mathcal{L}_{n}\;=\;-\lambda\,\frac{c^{n+1}-c^{n}}{\Delta t\,c^{n}}\,R_{c}(c^{n})H_{c}(v^{n})\;+\;\frac{\gamma_{c}\beta}{2\Delta t}\,(\kappa^{n})^{2}.$
@@
$\mathcal{J}[\kappa]\;=\;\int_{0}^{T}\Big[-\lambda\,\frac{\dot{c}(t)}{c(t)}\,R_{c}(c(t))H_{c}(v(t))\;+\;\frac{\tilde{\gamma}_{c}\beta}{2}\,\kappa(t)^{2}\Big]\,\mathrm{d}t.$ (6)
@@
In order to pass to a mean-field description of contacts, we suppose that the process of network formation is the result of small increments of the number of connections of the agents. This translates in the scaling of equation (3). Let $\varepsilon>0$, the scaled equation now reads
$\color[rgb]{0,0,0}{c^{\prime}=c-\left(\Psi^{\varepsilon}\left(\frac{c}{\overline{c}}\right)+\varepsilon\beta\left(\Phi(v)-\kappa_{\ell}\right)\right)c+\eta_{\varepsilon}c}$}
where $\eta_{\varepsilon}$ is a random variable with mean equal to zero, standard deviation $\sqrt{\varepsilon}\nu$, and $\mathbb{E}\left[\left(\eta_{\varepsilon}\right)^{3}\right]=O(\varepsilon^{3/2})$, while the value function $\Psi^{\varepsilon}$ is defined as
$\color[rgb]{0,0,0}{\Psi^{\varepsilon}(s)=\left(\frac{\mu}{1-\mu}\right)\frac{s^{\varepsilon}-1}{\frac{1+\mu}{1-\mu}s^{\varepsilon}+1}}$},
@@
Supposing the opinion variable $v$ fixed, we can now derive an evolutionary model for the controlled formation of the network conditioned on a given opinion.
To this end, we introduce the scaled conditional density of contacts defined through the decomposition $f_{\ell,\varepsilon}(v,c,t)=h_{\ell,\varepsilon}(c,t|v)g_{\ell,\varepsilon}(v,t),$ where $g_{\ell,\varepsilon}(v,t)$ is the scaled opinion marginal and $h_{\ell,\varepsilon}(c,t|v)$ denotes the scaled conditional distribution of contacts among agents having opinion $v$.
In this conditional description the opinion variable acts as a “frozen” parameter during the contact interaction step, while the kinetic evolution is performed only with respect to the contact variable $c$. Since the contact interaction does not modify the opinion variable, the opinion marginal $g_{\ell,\varepsilon}(v,t)$ remains unchanged during this interaction step, allowing us to write a closed kinetic equation for the conditional density $h_{\ell,\varepsilon}(c,t|v)$, which obeys a Boltzmann-like equation of the form
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}_{+}}\varphi(c){\color[rgb]{0,0,0}{h_{\ell,\varepsilon}(c,t|v)}}\mathrm{d}c=\mathbb{E}\left[\int_{\mathbb{R}_{+}}\chi(c)\left(\varphi(c^{\prime})-\varphi(c)\right){\color[rgb]{0,0,0}{h_{\ell,\varepsilon}(c,t|v)}}\mathrm{d}c\right],$ (7)
@@
$\color[rgb]{0,0,0}{\mathbb{E}\left[c^{\prime}-c\right]=-\Psi^{\varepsilon}\left(\frac{c}{\overline{c}}\right)c-\varepsilon\beta\left(\Phi(v)-\kappa_{\ell}\right)c}$}, (8)
@@
$\color[rgb]{0,0,0}{\mathbb{E}\left[(c^{\prime}-c)^{2}\right]=\varepsilon^{2}\left(-\frac{1}{\varepsilon}\Psi^{\varepsilon}\left(\frac{c}{\overline{c}}\right)-\beta\left(\Phi(v)-\kappa_{\ell}\right)\right)^{2}c^{2}+\varepsilon\nu^{2}c^{2}}$}, (9)
@@
${\color[rgb]{0,0,0}{\mathbb{E}\left[(c^{\prime}-c)^{3}\right]=}}$ ${\color[rgb]{0,0,0}{\varepsilon^{3}\left(-\frac{1}{\varepsilon}\Psi^{\varepsilon}\left(\frac{c}{\overline{c}}\right)-\beta\left(\Phi(v)-\kappa_{\ell}\right)\right)^{3}c^{3}+\mathbb{E}[\eta_{\varepsilon}^{3}]c^{3}}}$ (10)
${\color[rgb]{0,0,0}{+3\varepsilon^{2}\nu^{2}\left(-\frac{1}{\varepsilon}\Psi^{\varepsilon}\left(\frac{c}{\overline{c}}\right)-\beta\left(\Phi(v)-\kappa_{\ell}\right)\right)c^{3}}}.$